### v1

@@ 5.4 Real-Application Experiments
We finally evaluate ZapRAID when it hosts real applications. To host Linux applications, we export our prototype as an NVMe-over-Fabrics (NVMe-oF) target with TCP transport [9] and deploy the NVMe-oF host in the same testbed machine, so that our ZapRAID prototype serves as a Linux block device. We allocate the block device with a logical space capacity of 1 TiB, format the block device as an EXT4 file system, and run different types of applications atop the file system. As in §5.3, we consider the four settings of open segments. Note that the amount of write traffic may exceed the logical space capacity. Thus, we reserve an additional 25% of the physical space capacity for garbage collection.
@@
We evaluate ZapRAID in large file writes using Sysbench (v1.0.20) [11] to generate file write patterns. We first generate 128 800-MiB files (i.e., 100 GiB of data), and then use Sysbench to continuously issue 64-KiB write requests to the files for 10 minutes. We use one thread and issue writes in asynchronous mode with direct I/O. We issue fsync at the end of each experimental run.
Figure 17 shows the write throughput results. For a single open segment (Figure 17(a)), the performance findings are similar to those in Exp#1, meaning that the file write workload can also benefit from the use of intra-zone parallelism in ZapRAID. Compared with ZoneWrite-Only, ZapRAID achieves 76.9% and 15.3% higher write throughput for 4-KiB and 16-KiB chunks, respectively. ZapRAID also achieves similar write throughput as ZoneAppend-Only. For multiple open segments (Figure 17(b)), all three schemes show similar write throughput for $(N_{s},N_{l})=(2,2)$ and ZapRAID shows 16.3% higher throughput than ZoneAppend-Only for $(N_{s},N_{l})=(1,3)$; the findings are similar to those in Exp#7 under large writes (Figure 12(c)).
@@
Figure 17 depicts the throughput, in the number of transactions per second (TPS). For a single open segment with 4-KiB chunks (Figure 17(a)), ZapRAID shows 16.0% higher write throughput than ZoneWrite-Only and has similar throughput compared with ZoneAppend-Only. The improvement over ZoneWrite-Only is limited as the workload does not generate a sufficient number of concurrent write requests. For other settings of open segments, all three schemes have similar performance.
 Refer to caption Refer to caption
 (a) Single segment (b) Multiple segments
 Refer to caption Refer to caption

### v2

@@ 5.4 Real-Application Experiments
We finally evaluate ZapRAID when it hosts real applications and compare it with ZoneWrite-Only and ZoneAppend-Only (note that similar application benchmarks are also found in [45]). Recall that ZapRAID achieves a high write performance gain, especially in write-intensive workloads dominated by small writes (e.g., by up to 77.2% higher write throughput than ZoneWrite-Only from Exp#1), while maintaining read performance (from Exp#2). In real application workloads mixed with reads and writes, ZapRAID may not exhibit the same significant write-only performance gain as in Exp#1, yet its performance improvement remains observable.
To host Linux applications, we export our ZapRAID prototype as an NVMe-over-Fabrics (NVMe-oF) target with TCP transport [10] and deploy the NVMe-oF host in the same testbed machine, so that the prototype serves as a Linux block device. We allocate the block device with a logical space capacity of 1 TiB, format the block device as an EXT4 file system, and run different types of applications atop the file system. As in §5.3, we consider the four settings of open segments. Note that the amount of write traffic may exceed the logical space capacity. Thus, we reserve an additional 25% of the physical space capacity for garbage collection.
@@
We evaluate ZapRAID in file writes using Sysbench (v1.0.20) [12] to generate file write patterns. We first generate 128 files of 800 MiB each (i.e., totaling 100 GiB), and then use Sysbench to issue continuous 4-KiB write requests to the files for 10 minutes. We use one thread to issue writes in asynchronous mode with direct I/O. We issue fsync at the end of each run.
Figure 18 shows the write throughput results. For a single open segment (Figure 18(a)), ZapRAID benefits from intra-zone parallelism, achieving 38.6% higher write throughput for 4-KiB chunks than ZoneWrite-Only, and similar write throughput for 16-KiB chunks. ZapRAID also matches the write throughput of ZoneAppend-Only. The write throughput of the three schemes and the performance gains of ZapRAID are not as high as in Exp#1, as the workload contains frequent small writes, which necessitate frequent file system metadata updates with high synchronization overhead and limited write concurrency. For multiple open segments (Figure 18(b)), all three schemes show similar write throughput for $(N_{s},N_{l})=(2,2)$. ZapRAID achieves 20.2% higher throughput than ZoneWrite-Only for $(N_{s},N_{l})=(1,3)$; the findings are similar to those in Exp#7 under small writes (Figure 12(a)), albeit with a lower improvement.
@@
Figure 18 depicts the throughput in transactions per second (TPS). For a single open segment with 4-KiB chunks (Figure 18(a)), ZapRAID shows 16.0% higher write throughput than ZoneWrite-Only and similar throughput to ZoneAppend-Only. The limited improvement over ZoneWrite-Only is due to the low number of concurrent write requests and around 55% of all requests are reads, which are not optimized by Zone Append. For other settings of open segments, all three schemes have similar performance.
[Uncaptioned image] [Uncaptioned image]
@@
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
@@
Exp#13 (Key-value stores).
We evaluate ZapRAID when it hosts RocksDB (v9.7.2) [7] and use the benchmarking tool db_bench [5] to issue key-value pairs; RocksDB is an LSM-tree-based [63] key-value store that issues large sequential I/Os. We focus on two workloads: fillrandom (i.e., filling an empty database with randomly ordered keys) and overwrite (i.e., randomly overwriting an existing database). We set the total number of key-value pairs as 12.5 million with a value size of 8,000 bytes. We disable compression, and use eight threads to issue requests. For each run, we reset the volume, issue the fillrandom workload, and finally issue the overwrite workload.
Figures 20 and 20 show the throughput in number of operations per second (KOPS). For the fillrandom and overwrite workloads under a single open segment with 4-KiB chunks (Figures 20(a) and 20(a)), ZapRAID achieves 22.4% and 24.3% higher throughput than ZoneWrite-Only, respectively. For other settings, the throughput of all three schemes is similar. The improvement of ZapRAID is less pronounced than in microbenchmarks (Exp#1 and Exp#7), as both fillrandom and overwrite workloads have significant read traffic (about 50% of all I/O traffic) from frequent LSM-tree compaction operations, which involve retrieving and rewriting key-value pairs; Zone Append does not optimize reads.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 19: Exp#13 (Key-value stores). Throughput of the fillrandom workload.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 20: Exp#13 (Key-value stores). Throughput of the overwrite workload.
Exp#14 (Concurrent applications).
We combine two types of write-intensive workloads with different write patterns to evaluate ZapRAID. Specifically, we use Sysbench [12] to generate file write patterns with small writes. We also host MySQL and run the oltp_write_only benchmark, which issues large writes with less than 10% of reads. Our setting mimics a virtualized environment that hosts multiple workloads simultaneously in an SSD RAID array [30]. For small file writes, we follow the same approach as in Exp#11. For the oltp_write_only benchmark, we use Sysbench to fill the MySQL database with 8 tables and 10 million rows, and then run DELETE, INSERT, and UPDATE queries with 64 threads. Before each run, we clear and reformat the block device as an ext4 file system, and then execute both benchmarks simultaneously for 10 minutes.
Figure 22 shows the write throughput results for file small writes. The findings are similar to Exp#11. ZapRAID achieves 26.3% and 34.5% higher write throughput under a single open segment with 4-KiB chunks (Figure 22(a)) and four open segments with $(N_{s},N_{l})=(1,3)$ (Figure 22(b)), respectively. Figure 22 depicts the throughput of the oltp_write_only benchmark in transactions per second (TPS). For a single open segment with 4-KiB chunks (Figure 22(a)), ZapRAID achieves 32.9% higher write throughput than ZoneWrite-Only and has similar throughput to ZoneAppend-Only. For other settings of open segments, all three schemes show similar performance.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 21: Exp#14 (Concurrent applications). Throughput of small file writes.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 22: Exp#14 (Concurrent applications). Throughput of the oltp_write_only benchmark.